OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -13, -8, -3, 2, ...

OpenStudy (anonymous):

Tn = (-13) + (n-1)(5) Tn = -13+5n

OpenStudy (anonymous):

a_n = -13 + 5(n - 1) a_n = -13 + 5(n + 2) a_n = -13 + 5(n + 1)

OpenStudy (anonymous):

I get that it's +5 but how do I figure out the next part where its (n + or - something)

OpenStudy (anonymous):

Oh okay! Thank you! Nevermind! :)

OpenStudy (anonymous):

oops sorry for the miscalculation i my second step it should be -18+5n It is a formula you could remember, it is a + (n-1) d . where a is first term and d is the difference between two terms

OpenStudy (anonymous):

Thank you!